OpenStudy (anonymous):

solve and show work 3.8(s-6)+4.6=-22 s=?

OpenStudy (anonymous):

3.8s - (3.8)(6) + 4.6 = -22 3.8s - 22.8 + 4.6 = -22 3.8s - 18.2 = -22 3.8s = -3.8 s = -1

OpenStudy (anonymous):

woot woot!:) <3 seashell

OpenStudy (anonymous):

crap she beat me to it :(

OpenStudy (anonymous):

:D